# Transitivity of the pure Hurwitz classes of quadratic post-critically finite polynomials

### v1

@@ 4. Algorithm
Step 4 Runs at most thrice. If $\Gamma_{1}$ has two components and $\Gamma$ has one component, Function 1 may be called at most twice. Then Function 8 or 9 may be called. In all other cases, Step 4 runs once.
Step 4 runs at most seven times. We consider the upper bound for each component, and then multiply it by two to consider that it runs for two components. If $C_{j}^{i}$ is strictly pre-periodic then we run the step once. If $C_{j}^{i}$ is in a cycle, then we may need to run Step 4 to call Function 1 a maximum of two times before we are able to apply run Step 4 in order to call Function 11. Therefore Step 4 runs at most three times per component. Step 4 runs a final time to verify that the portrait is of the correct form. Therefore Step 2 must run a maximum of seven times for the two components.
Step 4: runs at most eight times. The general structure of Step 4 is as follows: if $\Gamma^{i}$ and $\Gamma$ have different pre-period structures, first run Step 4 at most three times to obtain branched portrait with one pre-period. Then run Step 4 at most four more times to obtain the target pre-period structure (of $\Gamma$). Then run Step 4 a final time to verify that the input portrait has the same pre-period structure of $\Gamma$ and move to the next step. The maximum occurs if $\Gamma^{i}$ has no pre-periods and $\Gamma$ either has two disjoint pre-periods or distinct and intersecting pre-periods.
Step 4: Let $(\Gamma^{i},f_{i})$ be the pair returned from Step 4 or 4 (whichever called Step 4). Let $r_{(1,2)}^{i},r_{1}^{i},r_{2}^{i},r_{(1,2)},r_{1},$ and $r_{2}$ be as defined in Step 4. Then Step 4 will be called at most $|r_{(1,2)}^{i}-r_{(1,2)}|+|r_{1}^{i}-r_{1}|+|r_{2}^{i}-r_{2}|+1$ times in order to adjust the length of the intersection of the pre-periods of $C_{1}^{i}$ with $C_{2}^{i}$, the length of the pre-period of $C_{1}^{i}$ which is unique from $C_{2}^{i}$, and the length of the pre-period of $C_{2}^{i}$ which is unique from the pre-period of $C_{1}^{i}$ since all three lengths may be changed. Step 4 will run one remaining time, to verify that the pre-period lengths are all equal.
Therefore, Step 4 runs a maximum of $|r_{(1,2)}^{i}-r_{(1,2)}|+|r_{1}^{i}-r_{1}|+|r_{2}^{i}-r_{2}|+1$ times.
Step 4: Let $(\Gamma_{i},f_{i})$ be the pair returned from Step 4, 4, or 4 (whichever called Step 4). The number of times we call Step 4 depends on the number of cycles (hence the number of components), the number of cycle distances that we measure (either the cycle length or two lengths measured between critical points or pre-periods), and the difference of those lengths between $\Gamma^{i}$ and $\Gamma$.
Suppose that $\Gamma^{i}$ and $\Gamma$ both have two nontrivial portrait components. As in Step 4, let $k_{j}^{i}$ denote the length of the cycle of the component $\Gamma_{j}^{i}$, and let $k_{j}$ denote the length of the component $\Gamma_{j}$. We note that $k_{1}^{i}$, $k_{2}^{i}$, $k_{1}$ and $k_{2}$ must all be finite because $\Gamma^{i}$ and $\Gamma$ are both finite graphs. Then Step 4 runs at most $|k_{1}^{i}-k_{1}|+|k_{2}^{i}-k_{2}|+1$ times, where the final iteration is to verify that the $k_{1}^{i}=k_{1}$ and $k_{2}^{i}=k_{2}$.
Suppose now that $\Gamma^{i}$ and $\Gamma$ both have one portrait component, that neither critical point is periodic, and that the pre-periods of the critical points in $\Gamma^{i}$ (and hence also in $\Gamma$) intersect non-trivally. As in Step 4, let $k^{i}$ denote the length of the cycle of $\Gamma_{i}$ and let $k$ denote the length of the cycle of $\Gamma$. Observe that $k^{i}$ and $k$ are both finite because $\Gamma^{i}$ and $\Gamma$ are both finite graphs. Then Step 4 runs at most $|k^{i}-k|+1$ times, where the final iteration is to verify that $k^{i}=k$.
Suppose now that $\Gamma^{i}$ and $\Gamma$ both have one portrait component, and that $\Gamma^{i}$ and $\Gamma$ both have no pre-periods (two periodic critical points), one pre-period (one periodic critical point and one strictly pre-periodic critical point), or two disjoint pre-periods. Let $k_{1}^{i},k_{2}^{i},k_{1},$ and $k_{2}$ be defined as in Step 4. We note that $k_{1}^{i}$, $k_{2}^{i}$, $k_{1}$ and $k_{2}$ must all be finite because $\Gamma^{i}$ and $\Gamma$ are both finite graphs. If $\Gamma$ has one pre-period and $C_{2}$ is the first periodic point of the cycle, then Step 4 runs $k_{1}^{i}+|(k_{1}^{i}+k_{2}^{i})-k_{2}|+1$ times. Otherwise Step 4 runs a total of $|k_{1}^{i}-k_{1}|+|k_{2}^{i}-k_{2}|+1$ times, where the final iteration verifies that the lengths are equal and terminates the algorithm.
From this analysis, we see that Step 4 must run a finite number of times.
@@
The final ingredient is a more general restatement of [7, Lemma 3.6]. Let $Q(n)$ be the set of quadratic branched covers with $n$ post-critical points. The proof is identical to that of Kelsey–Lodge, by replacing the set $Q(4)^{*}$ with $Q(n)$.
@@
Let $f,g\in Q(n)$. Then $f,g$ are pure Hurwitz equivalent if and only if they have equivalent portraits.
@@
Let $f$ and $g$ be the branched covers indicated in the theorem. Let $\Gamma_{f}$ be the portrait of $f$ and let $M_{f}$ be the marked set of $f$. Let $\Gamma_{g}$ be the portrait of $g$. We run the algorithm by inputting $(\Gamma_{f},f)$, $\Gamma=\Gamma_{g},$ the marked set $M_{f}$, the identity mapping class, and the sequence $\{(\Gamma_{f}=\Gamma_{0},f=f_{0})\}$ into Step 4. Let $(\Gamma_{n},f_{n},M_{n},\phi_{n},\{(\Gamma^{0},f_{0}),(\Gamma^{1},f_{1}),\cdots,(\Gamma^{i},f_{i})\})$ be the return of Step 4 when the cycle terminates. By construction $\phi_{i}\in\operatorname{Mod}(S^{2},M_{i})$ for all $0\leq i\leq n$. In particular, $\phi_{n}\in\operatorname{Mod}(S^{2},\phi_{n})$ is the mapping class desired in the statement of the theorem. Next we prove that $f\circ\phi_{n}$ has the same portrait as $g$.
Let $i_{1}$ be such that $\Gamma^{i_{1}}$ is the portrait output by the final iteration of Step 4, that is, the initial input into Step 4 or 4. Then $\Gamma_{i_{1}}$ will have the same number of components as $\Gamma$.
Step 4 calls Step 4 or 4. Because Step 4 is only called by Step 4 and itself and Step 4 is only called by Step 4 and itself, only one of Step 4 or Step 4 will run (not both). Let $i_{2}$ be such that $(\Gamma^{i_{2}},f_{i_{2}})$ is the pair output by the final iteration of either Step 4 or Step 4 (whichever one was called), that is, the initial input into Step 4. Let $i_{3}$ be such that the portrait $\Gamma^{i_{3}}$ output by the final iteration of Step 4, Step 4, or Step 4 (whichever calls Step 4), that is, the first input into Step 4. Because Steps 4, 4, and 4 all may call Step 4, it is possible that $i_{2}=i_{3}$.
We claim that:
  1. (1)
    for all $i\geq i_{1}$, $\Gamma^{i}$ and $\Gamma$ have the same number of (non-empty) components,
  2. (2)
    for all $i\geq i_{2}$, $\Gamma^{i}$ and $\Gamma$ have the same pre-period structure,
  3. (3)
    for all $i\geq i_{3}$, all pre-periods lengths of $\Gamma^{i}$ are the same as the lengths of the corresponding pre-periods in $\Gamma$, and
  4. (4)
    the final portrait $\Gamma^{n}$ is isomorphic to the portrait $\Gamma$.
First we observe that $\Gamma^{i_{1}}$ and $\Gamma$ have the same number of components. Steps 4-4 use Functions 1-7 and 10-13, none of which change the number of components of $\Gamma^{i}$. Thus the first claim is proved.
For the second claim, observe that the final iteration of both Steps 4 and 4 verifies that $\Gamma^{i_{2}}$ and $\Gamma$ have the same pre-period structure before moving to Step 4 and 4. Step 4 only uses Functions 3 and 4. Step 4 calls Functions 1, 2, 5(E) and 5(L). None of these functions change the pre-period structure of $\Gamma^{i}$. Therefore the second claim is true.
To verify the third claim, observe that final iteration of Step 4 ensures that the lengths of the pre-periods in $\Gamma^{i_{3}}$ have the same lengths as the corresponding pre-periods in $\Gamma$. To verify the third claim, we observe that Step 4 calls Functions 1, 2, 5(E) and 5(L). None of these functions change the lengths of pre-periods of $\Gamma^{i}$ for $i\geq i_{3}$. Therefore the third claim is true.
The final iteration of Step 4 ensures that the lengths within the cycle(s) of $\Gamma^{n}$ are the same as the corresponding lengths in $\Gamma$. Then by Proposition 2.1, the portrait $\Gamma^{n}$ is isomorphic to $\Gamma$. Thus the claims are verified.
Finally, because the portraits of $f_{n}$ and $g$ are equivalent, by Lemma 5.2 (cf. [7, Lemma 3.6]), $f_{n}$ and $g$ are pure Hurwitz equivalent. ∎

### v2

@@ 4. Algorithm
Step 4 Runs twice: once for $g$ and once for $h$.
Step 4 runs at most six times. The step runs at most once for each component of both $g$ and $h$, plus a final time (for each $g$ and $h$) to verify that both components are cycles: a total of six times.
Step 4: We consider the number of times Step 4 is run for both $g$ and $h$. Let $(\Gamma^{i},f_{i})$ be the pair returned from Step 4 when the algorithm is run for $g$ and let $(\Delta^{i},f_{i})$ be the pair returned from step 4 when the algorithm is run for $h$. Let $k_{j}^{g}$ be the cycle length of $\Gamma^{i}_{j}$ (when the algorithm is run for $g$) and $k_{j}^{h}$ be the cycle length of $\Delta_{i}^{j}$ (when the algorithm is run for $h$). Then Step 4 runs $k_{1}^{g}+k_{2}^{g}-1$ for $g$. Indeed, Step 4 runs $k_{1}^{g}-1$ times for the first component, $k_{2}^{g}-1$ times for the second component and one additional time to check that the lengths are correct. A similar calculation shows that Step 4 runs $k_{1}^{h}+k_{2}^{h}-1$ for $h$. Hence Step 4 runs a total of $k_{1}^{g}+k_{2}^{g}+k_{1}^{h}+k_{2}^{h}-2$ times.
@@
The final ingredient is a more general restatement of [8, Lemma 3.6]. Let $Q(n)$ be the set of quadratic branched covers with $n$ post-critical points. The proof is identical to that of Kelsey–Lodge, by replacing the set $Q(4)^{*}$ with $Q(n)$.
@@
Let $f,g\in Q(n)$. Then $g,h$ are pure Hurwitz equivalent if and only if they have equivalent portraits.
@@
Let $g$ and $h$ be the branched covers indicated in the theorem. Let $\Gamma_{g}$ be the portrait of $g$ and let $M_{g}$ be the marked set of $g$. Let $\Gamma$ be the portrait of $z^{2}$. We first run the algorithm by inputting $(\Gamma_{g},g)$, the marked set $M_{g}$, the identity mapping class, and the sequence $\{(\Gamma_{g}=\Gamma_{0},g=f_{0})\}$ into Step 4. Let $(\Gamma_{n},f_{n},M_{n}^{f},\phi_{n}^{f},\{(\Gamma^{0},f_{0}),(\Gamma^{1},f_{1}),\cdots,(\Gamma^{n},f_{n})\})$ be the return of Step 4 when the cycle terminates. By construction $\phi_{i}^{g}\in\operatorname{Mod}(S^{2},M_{i}^{g})$ for all $0\leq i\leq n$. In particular, $\phi_{n}^{g}\in\operatorname{Mod}(S^{2},M_{n}^{g})$. Next we prove that $g\circ\phi^{f}_{n}$ has the same portrait as $z^{2}$.
First we observe that $\Gamma^{1}$ and $\Gamma$ both have two components by applying Step 4. Steps 4 and 4 call Functions2 and 3, neither of which changes the number of components of $\Gamma^{i}$.
Next, let $i_{2}$ be the index of the final iteration of Step 4 (that calls Step 4). Observe that $\Gamma^{i_{2}}$ has two components that are each cycles, just like $\Gamma$. The final iteration of Step 4 verifies that $\Gamma^{i_{2}}$ and $\Gamma$ have the same pre-period structure before moving to Step 4. Step 4 calls Function 3, which does not change pre-period structure of $\Gamma^{i}$.
The final iteration of Step 4 ensures that $\Gamma^{n}_{f}$ consists of two components, both of which are cycles of length 1. This is the same portrait as $\Gamma$.
We repeat the process for $h$ by inputting $(\Gamma_{h},h)$, the marked set $M_{h}$, the identity mapping class and the sequence $\{(\Gamma_{h}=\Gamma_{0},h=f_{0})\}$ into Step 4. The final iteration of Step 4 will output a sequence $(\Gamma_{m},f_{m},M_{m}^{h},\phi_{m}^{h},\{(\Gamma^{0},f_{0}),(\Gamma^{1},f_{1}),\cdots,(\Gamma^{m},f_{m})\})$. Similarly to $g$, we we obtain the mapping class $\phi_{m}^{h}\in\operatorname{Mod}(S^{2},M_{m}^{h})$. The final portrait is the same as $\Gamma$.
Then $g\circ\phi_{n}^{g}$ and $h\circ\phi_{m}^{h}$ both have portraits isomorphic to $\Gamma$. Let $\phi=\phi_{n}^{g}\circ(\phi_{m}^{h})^{-1}\in\operatorname{Mod}(S^{2},M_{n}^{g}\cup M_{m}^{h})$. Then $g\circ\phi=(g\circ\phi_{n}^{g})\circ(\phi_{m}^{h})^{-1}$ has portrait isomorphic to the portrait of $h$.
Finally, because the portraits of $g\circ\phi$ and $h$ are equivalent, by Lemma 5.2 (cf. [8, Lemma 3.6]), $g\circ\phi$ and $h$ are pure Hurwitz equivalent. ∎